OpenStudy (anonymous):

what is 3/15+12/18 in lowest terms

OpenStudy (anonymous):

13/15

OpenStudy (anonymous):

do u want me to draw it ? i plugged it into the calculator but there is a way to do it by hand

OpenStudy (anonymous):

(1/5)+(2/3)

OpenStudy (anonymous):

=(3+10)/15=13/15

OpenStudy (anonymous):

in lowest terms

OpenStudy (anonymous):

(13/15) is the lowest terms it cant be divided by anything without getting another fraction

OpenStudy (anonymous):

no 3/15+12/18 in lowest terms

OpenStudy (phi):

that is what people posted up above they first said: 3/15 can be reduced to 1/5 (divide top and bottom by 3) reduce 12/18 to 2/3 (divide top and bottom by 6) add 1/5 + 2/3 common denominator is 15 can you finish?

OpenStudy (anonymous):

not really i hate fraction lol could u help me with afew please

OpenStudy (phi):

In fact, watch the first few (they are all short) http://www.khanacademy.org/math/arithmetic/fractions

OpenStudy (anonymous):

what is the lcm of 6 and 9

OpenStudy (phi):

write 6 as 2*3 write 9 as 3*3 you need 1 two and 2 3's so 2*3*3= 18

OpenStudy (anonymous):

and thank you u was alot of help lol